OpenStudy (anonymous):

square root of 30,000 simplify!

OpenStudy (anonymous):

Google it

OpenStudy (icecreamyummy):

you can not square it I used mathway.com

OpenStudy (freckles):

\[\text{ hint } \sqrt{30000}=\sqrt{3 \cdot 10000}=\sqrt{3 \cdot 100^2}\]

OpenStudy (triciaal):

30000 = 3.0 * 10^4 sqrt of 30000 = sqrt (3 * 10^4) = sqrt 3 * sqrt 10^4 = sqrt 3 * 10^2 1.732 *100 = 173.2

OpenStudy (triciaal):

@ICECREAMyummy the question ask for the square root if you wanted to find the square you can also use scientific notation to simplify

OpenStudy (triciaal):

@ICECREAMyummy 30000^2 would be 3*10^4raised to power 2 =3^2*10^4^2 = 9*10^8

OpenStudy (icecreamyummy):

I am I 8TH grade soo @triciaal

OpenStudy (triciaal):

@ICECREAMyummy what does that mean?

OpenStudy (icecreamyummy):

it means that I have not leard the square roots

OpenStudy (triciaal):

ok the opposite of square is square root like multiplication is the opposite of division addition is the opposite of subtraction you also need to read the question properly it will make a difference later.

OpenStudy (triciaal):

example 6 square is 36 so the square root of 36 is 6 means what number when multiplied by itself = 36

OpenStudy (triciaal):

you should be ready to lead the class now